OpenStudy (anonymous):

\[\int\limits_{2\pi}^{3\pi} \left| sinx/2 \right|dx\]

OpenStudy (experimentx):

split the limit make function such that sine is always positive in 2pi to 3pi

OpenStudy (experimentx):

the best start would be to draw graph first.

OpenStudy (anonymous):

split it as in \[\int\limits_{2\pi}^{3\pi}\left| sinx \right|dx + \int\limits_{2\pi}^{3\pi}\left| 1/2 \right|dx\]

OpenStudy (experimentx):

so your function was 1/2 * sin(x)

OpenStudy (anonymous):

no my function was the absolute value of sinx/2

OpenStudy (anonymous):

oh so rather than the + sign it should be a - sign

OpenStudy (experimentx):

was it sin(x/2) for sin(x)/2 ??

OpenStudy (anonymous):

it's sinx/2

OpenStudy (experimentx):

can put plot that function in wolframalpha ??

OpenStudy (anonymous):

i don't know how to use wolfram

OpenStudy (anonymous):

im not good with graphing :(

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=plot+ |sinx%2F2|from+x%3D2pi+to+3pi

OpenStudy (experimentx):

and make sure if 1/2 in inside sin or outside?? http://www.wolframalpha.com/input/?i=graph+sin%28x%2F2%29+from+3pi+to+4pi

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=graph+ |sin%28x%2F2%29|+from+3pi+to+4pi

OpenStudy (experimentx):

it's negative so ... integrate -sinx instead of |sinx| ... if you know how to integrate it should be all right.

OpenStudy (anonymous):

im so confused

OpenStudy (experimentx):

you know how to integrate right??

OpenStudy (anonymous):

can u explain to me step wise on how to solve it. i'm studying for a test and i don't know how to solve problems like these

OpenStudy (anonymous):

a little.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integrate+ |sin%28x%29%2F2|+from+2pi+to+3pi

OpenStudy (experimentx):

can you integrate 1/2*sin(x)

OpenStudy (anonymous):

i don't know how to integrate with |absolute value|

OpenStudy (anonymous):

i know that we can take the constant 1/2 out so it can written as \[1/2\int\limits_{2\pi}^{3\pi}|sinx| dx\]

OpenStudy (experimentx):

in place of |sin x| you can place - sin x because sin x will be negative in interval 2pi to 3pi

OpenStudy (anonymous):

\[\int\limits_{}^{} sinx dx= cosx \] i know that

OpenStudy (anonymous):

i meant -cosx sorry

OpenStudy (experimentx):

\(\huge \int_{2\pi}^{3\pi}|\sin x| dx = \int_{2\pi}^{3\pi}-\sin x dx \)

OpenStudy (anonymous):

ok. so when i integrate that i'll get positive cosx

OpenStudy (experimentx):

|x| = -x or |x| = +x you can change values according to make x always positive. YES

OpenStudy (anonymous):

so after integrating it's 1/2cos(3pi)-1/2cos(2pi)

OpenStudy (experimentx):

yes ...thats your answer. Sorry about the limit, .... the limit was mistake.

OpenStudy (anonymous):

ok thanks you!!!

OpenStudy (experimentx):

yw .. np